Grade 10Mechanics

Find the ratio of a systolic blood pressure of 120 mm of Hg to atmospheric pressure . Standard atmospheric pressure is 1.01 X 105 Pa.

Aiden Mathew Denny
8 Years agoGrade 10

1 Answer

Venkat
ApprovedApproved Tutor Answer8 Years ago
Pressure due to the 120 mm of mercury =  ρgh = (13,600 kg/m3)(9.8 m/s2)(0.120 m) =  0.16 x 10Pascals.
 
Ratio = 0.16/1.01 = 0.16